A trunk with a weight of 210 N rests on the floor. The coefficient of static friction between the trunk and the floor is 0.40, while the coefficient of kinetic friction is 0.29. What is the minimum magnitude for a horizontal force with which a person must push on the trunk to start it moving?
1
Expert's answer
2011-10-20T08:30:45-0400
There are two forms of friction, kinetic and static. If you try to slide two objects past each other, a small amount of force will result in
no motion. The force of friction is greater than the applied force. This is
static friction. If you apply a little more force, the object "breaks free" and
slides, although you still need to apply force to keep the object sliding. This
is kinetic friction. You do not need to apply quite as much force to keep the
object sliding as you needed to originally break free of static
friction.

Magnitude for a horizontal force: F=P*max(0,40;0.29)=210*0.4=84N